OpenStudy (anonymous):

reduce to lowest terms 3m^2-3n^2/6m-6n

OpenStudy (anonymous):

\[\frac{3m^2-3n^2}{6m-6n}=\frac{3(m+n)(m-n)}{6(m-n)}=\frac{m+n}{2}\]